# Effective and Efficient Cross-City Traffic Knowledge Transfer: A Privacy-Preserving Perspective

### v1

@@ 4.5. Training Process
Training Algorithm. For convenient reproduction, we provide detailed training Algorithm 1 of the FedTT framework, including the client and server. In the client (line 1), it first predicts the missing traffic data (line 2). During each training round (line 3), it transforms the data from the traffic domain of the source city to that of the target city (line 4). Then, it gets the aggregated data of the last round from the server (line 5) and classifies the traffic domain of the data (line 6). Finally, it computes the mask data at the current round and sends it to the server (lines 7–8). In the server (line 9), during each training round (line 10), it gets the mask data from clients and computes the aggregated data at the current round (lines 11–12). Then, it classifies the traffic domain of the data (line 13). Next, it predicts future traffic data based on historical traffic observations (line 14). Finally, it sends the aggregated data to clients (line 15).
Complexity Analysis. We also give the complete complexity analysis for the training of the FedTT framework, i.e., Algorithm 1. For the client $c_{i}$, the training complexity is $O((|\mathcal{M}^{R_{i}}|+|\mathcal{M}^{S}|)\times|\mathcal{X}^{R_{i}}|)$ at each round. For the server $s$, the training complexity is $O((|\mathcal{M}^{S}|+\textit{MC}(\theta_{\textit{TP}}))\times(|\mathcal{X}^{S}|+\sum_{i=1}^{n}|\mathcal{X}^{R_{i}}|))$ at each round, where $\textit{MC}(\theta_{\textit{TP}})$ is the model complexity of the traffic model $\theta_{\textit{TP}}$.
@@ 5.1. Experimental Settings
We use four datasets to evaluate the proposed FedTT framework in experiments, which are widely used in traffic prediction tasks (Zhao et al., 2023; Ji et al., 2023; Jiang et al., 2023), as shown in Table 2. Specifically, PeMSD4 (P4(pem, 2024), PeMSD8 (P8(pem, 2024), FT-AED (FT(Coursey et al., 2024), and HK-Traffic (HK(hk, 2024) were collected in the San Francisco, San Bernardino, Nashville, and Hong Kong, respectively. Among them, three datasets are considered as three source cities, and one dataset serves as the target city, leading to four combinations: (P8, FT, HK) $\to$ P4, (P4, FT, HK) $\to$ P8, (P4, P8, HK) $\to$ FT, and (P4, P8, FT) $\to$ HK. Besides, we select traffic flow, speed, and occupancy prediction tasks for experiments, which are also widely studied in the community (Zhao et al., 2023; Ji et al., 2023; Jiang et al., 2023).

### v4

@@ 4.5. Training Process
Training Algorithm. For convenient method reproduction, we provide detailed training Algorithms 1 and 2 of the FedTT framework, including the client and server.
In the client (i.e., Algorithm 1), the target city acts as the server. Before the training process, the client completes the missing traffic data through the traffic view imputation method (line 1). During each training round and each traffic data (lines 2–3), it first transforms the data from the traffic domain of the source city to that of the target city using the local generator model (line 4) and classifies the transformed data using the local discriminator model (line 5). If the training process is in the first round using the first data instance (line 6), the client encrypts the transformed data using homomorphic encryption and sends it to the server (lines 7-8). Otherwise, if the training process is in the first round using the second data instance (lines 9-10), the client gets the encrypted data and decrypts it to get the previous aggregated data (lines 11-12). For subsequent rounds or data instance, the client directly gets the previous aggregated data from the server without decryption (lines 13-14). In either case, it classifies the previous aggregated data using its local discriminator model (line 15). Then it masks the transformed data using the previous aggregated and transformed data (line 16). Finally, it sends the mask data to the server for data aggregation (lines 17).
In the server (i.e., Algorithm 2), the source cities act as the clients. During each training round and each traffic data (lines 1–2), if the training process is in the first round using the first data instance (line 3), the server gets the encrypted data from clients (line 4). Then, it aggregates them by summing up, and send the aggregated encrypted data to back to the clients for further processing (lines 5-6). For subsequent rounds or data instances (line 7), the server gets the mask data from clients (line 8). Then, it aggregates the masked data using the previous aggregated data (line 9). Next, it classifies the aggregated data using its global discriminator model and sends the aggregated data back to the clients (lines 10–11). Finally, at the end of each training round, it classifies local traffic data and performs traffic prediction using the aggregated and local traffic data (lines 12–13).
Complexity Analysis. We also give the complete complexity analysis for the training of the FedTT framework, i.e., Algorithms 1 and 2. For the client (i.e., Algorithm 1), the training complexity is $O((|\mathcal{M}^{R_{i}}|+|\mathcal{M}^{S}|)\times(F_{1}\times H)^{2}\times|\mathcal{X}^{R_{i}}|)$ at each round. For the server (i.e., Algorithm 2), the training complexity is $O((|\mathcal{M}^{S}|\times(F_{1}\times H)^{2}+\textit{MC}(\theta_{\textit{TP}}))\times(|\mathcal{X}^{S}|+\sum_{i=1}^{n}|\mathcal{X}^{R_{i}}|))$ at each round. Here, $|\mathcal{M}^{R_{i}}|$ and $|\mathcal{M}^{S}|$ are the number of sensors in the source city $R_{i}$ and target city $S$, respectively. Besides, $|\mathcal{X}^{R_{i}}|$ and $|\mathcal{X}^{S}|$ are the number of traffic data in the source city $R_{i}$ and target city $S$, respectively. In addition, $F_{1}=3$ is the dimensions of traffic data features, and $H=1024$ is the hidden dimensions of the three-layer MLP model in $\theta_{\textit{Gen}}^{R_{i}}$ and $\theta_{\textit{Dis}}^{R_{i}}$. Moreover, $\textit{MC}(\theta_{\textit{TP}})$ is the model complexity of $\theta_{\textit{TP}}$ (i.e., $\theta_{\textit{DyHSL}}$).
@@ 5.1. Experimental Settings
We use four datasets to evaluate the proposed FedTT framework in experiments, which are widely used in traffic prediction tasks (Zhao et al., 2023; Ji et al., 2023; Jiang et al., 2023), as shown in Table 2. Specifically, PeMSD4 (P4(2), PeMSD8 (P8(2), FT-AED (FT(Coursey et al., 2024), and HK-Traffic (HK(53) were collected in the San Francisco, San Bernardino, Nashville, and Hong Kong, respectively. Among them, three datasets are considered as three source cities, and one dataset serves as the target city, leading to four combinations: (P8, FT, HK) $\to$ P4, (P4, FT, HK) $\to$ P8, (P4, P8, HK) $\to$ FT, and (P4, P8, FT) $\to$ HK. Besides, we select traffic flow, speed, and occupancy prediction tasks for experiments, which are also widely studied in the community (Zhao et al., 2023; Ji et al., 2023; Jiang et al., 2023).